### v8

# Design and Implementation of a Scalable Financial Exchange in the Public Cloud

Abstract
Financial exchanges are migrating to the cloud, but the best-effort nature of the public cloud is at odds with the stringent latency requirements of exchanges. We present Jasper, a system for meeting the networking requirements of financial exchanges on the public cloud. Jasper uses an overlay tree to scalably multicast market data from an exchange to ~1000 participants with low latency () and difference in data reception time between any two participants. Jasper reuses the same tree for scalable inbound communication (participants to exchange), augmenting it with order pacing and a new priority queue, Limit Order Queue (LOQ), to efficiently handle bursts of market orders. Jasper achieves better scalability and 50% lower latency than the AWS multicast service [1]. During bursty market activity, LOQ nearly doubles the order processing rate.
1 Introduction
Financial exchanges are migrating to the public cloud for several reasons such as improved scalability and reduced capital expenditure. Despite its benefits, the cloud poses unique challenges. In particular, exchanges have traditionally operated in on-premise or colocation facilities, engineered for deterministic and low latency. For fair market access, exchanges provide deterministic latency by equalizing cable lengths between the exchange and participant servers; thus, all participants receive market data simultaneously. To reduce latency, exchanges often use FPGAs as gateways [2] and L1 switches that perform low-latency physical-layer multicast [3].
However, the public cloud lacks these enhancements. It is a best-effort environment characterized by non-determinism (e.g., latency variance in Figure 1). In response, several prior projects have developed techniques for cloud-based exchanges. These include using synchronized clocks to compensate for non-determinism [4], using SmartNICs [5], and new definitions of fairness [6, 7]. These projects show promising results for 10s of participants. Our paper develops new networking techniques for scale: how do we architect an exchange to support communication between the exchange and ~1000 participants in the cloud? In practice, a scalable cloud exchange would also need scale-out compute techniques for the exchange server—a concern that is out of scope.
Our system, Jasper, tackles 3 challenges to provide network support for scalable exchanges. First, how do we support outbound communication of market data from the exchange to ~1000 market participants, while ensuring (a) low spatial variance, i.e., all participants receive market data nearly simultaneously, (b) low latency from the exchange server to the participant, and (c) low temporal variance, i.e., latency doesn’t fluctuate over time? Second, how do we support inbound communication of participant orders to the exchange—especially during intense market activity when bursts of orders from many participants arrive at once, causing incast-style [8] drops and increased order latency? Third, trust: a rational participant will act on market data as soon as it is available. Hence, for fairness, the network must hold data and release it to a given participant only when it is sure that the data is available to all participants. How do we mitigate this additional holding latency and latency variance? In tackling these challenges, we make 3 contributions.
First, to scale outbound communication from the exchange to a large number of participants, we employ an overlay multicast tree composed of a root exchange VM, proxy VMs as intermediate nodes, and participant VMs as leaf nodes. We develop a simple and effective heuristic to tune the tree’s fan-out and depth, navigating a tradeoff between increased serialization delay due to greater fan-out and increased propagation delay because of greater depth. To lower latency and counter the cloud’s variability, we pervasively employ hedging: routing redundant copies of the market data through multiple proxy and receiver VMs and rotating parent-to-child associations at each tree level on each multicasted packet.
Second, to scale inbound communication from participants to the exchange, we use the tree in reverse; this reduces the fan-in at the matching engine, decreasing incast-style drops. To further reduce drops during bursts of orders, every tree node paces inbound transmissions when the one-way delay between that node and its parent is high. While pacing reduces drops, it builds up queues. To counter this queuing latency, we develop a specialized scheduling algorithm, limit-order queuing (LOQ), which runs on the proxy VMs, and transmits market orders closest to the midpoint of the bid-ask spread—rather than transmitting FIFO. By prioritizing orders close to the midpoint, LOQ improves the likelihood that transmitted orders will be matched immediately. This improves matching throughput and order-matching latency relative to FIFO.
Third, we explore a range of trust models to capture how participants’ trading algorithms interface with the leaf nodes on the multicast tree. These include running the leaf within a VM and the algorithm within a protected enclave in the same VM, running the two entities as 2 separate cloud VMs, and running the 2 entities as 2 separate functions in a process. We describe and measure the performance vs. security tradeoffs inherent in each solution.
Jasper can support a maximum throughput of 175K multicast messages per second at which point it is limited by a proxy VM’s egress bandwidth. It also scales to 1000 receivers, achieving a median multicast latency of while achieving a spatial variance of across these receivers. Jasper also scales better and achieves 50% lower latency compared to AWS’ Transit Gateway-based multicast. On the inbound side, Jasper efficiently handles large bursts of orders, using LOQ to nearly double the order-matching rate and maintain low order latencies during bursts.
To contextualize Jasper’s absolute performance, Jasper falls short of heavily engineered on-premises financial exchanges, which achieve spatial variance of a few nanoseconds [2]. Nevertheless, for certain financial instruments (e.g., cryptocurrency), their dominant exchanges [9, 10] have been built on the cloud since the beginning. For such instruments, Jasper is very competitive. Further, participant-to-exchange interfaces exist on a usability-performance tradeoff curve: colocated participants at one end and web APIs at the other [11, 12]. Jasper presents a compelling point on this curve: low and predictable latencies of ~ while scaling to 1K market participants and many additional benefits of the cloud.
2 Background
Financial Exchange Setup. An exchange typically has an exchange server and multiple market participant (MP) servers. The exchange server runs a matching engine (ME) to process trading orders from the MPs and multicast market data to them. Market data reveals market state e.g., asset prices and processed orders. Orders can be bid orders, which aim to purchase an asset at a specific price, and ask orders, which aim to sell an asset at a specific price.
The ME maintains a limit order book (LOB) (Fig. 3), which lists all bid and ask orders from MPs. When a bid price exceeds the lowest ask price, the orders are matched and executed. Unexecuted orders remain in the LOB, waiting for a match. In the price-time-priority algorithm [13], orders are arranged in price levels, with those at the same level sorted by arrival time, prioritizing older orders.111Price-time-priority is widespread, so we use it. Alternatives exist such as the pro-rata algorithm [13] An LOB snapshot shows a separation between bid and ask price levels, with the mid-price indicating the asset’s true value. Thus, orders closer to the mid-price have higher chances of getting matched early. The mid-price changes with new orders.
Traditionally, the exchange and MP servers are colocated in the same datacenter and connected via equal-length wires using L1 [3] switches with switch multicast. The overall infrastructure is designed to provide low and equal latency from the exchange server to all MPs in both directions.
Challenges of Cloud Migration. While the public cloud offers many advantages, it does not offer low-level control, e.g., allowing tenants to control wire lengths. The public cloud also exhibits high latency variance [14] and compute variance [15]. The latency between a pair of VMs can be significantly different from another VM pair [16]. Latency also fluctuates over time: Figure 1 plots the latency between a pair of VMs in an AWS region for a tumbling window of . The figure also shows infrequent but unpredictable latency spikes that increase latency substantially [14, 7]. Such phenomena make it challenging to achieve low and deterministic latency.
Given the above challenges, it is reasonable to ask whether financial exchanges should ever be migrated to the cloud—and if so, how. This is an ongoing debate with reasonable arguments on both sides. Beyond the research projects [4, 7, 5] discussed below, some cloud providers are exploring close partnerships to build private clouds for exchanges’ bespoke requirements [17, 18]. Jasper contributes to this debate by empirically studying what performance guarantees can be achieved if a cloud tenant designs the exchange on the public cloud with publicly available cloud APIs. With Jasper’s DIY approach, a tenant does not need to wait for a cloud provider to build private clusters. While Jasper doesn’t achieve on-prem performance, its performance is competitive with several cloud-based exchanges [10, 9].
Prior Work and Jasper’s Motivation. CloudEx [4] synchronizes the clock of the exchange server VM to the clocks of several gateway VMs, to which participant VMs connect. On the outbound side, the exchange sets a release time for market data which the gateways respect to provide simultaneous delivery of market data. On the inbound side, each gateway timestamps its incoming participant’s orders; this timestamp determines an order’s priority in the price-time-priority algorithm. Octopus [5] implements an approach similar to CloudEx using exchange-controlled SmartNICs on the participant’s machines instead of gateway VMs to implement simultaneous release of market data. Recent literature including DBO [7], FBA [19], and Libra [20], proposes new definitions of fairness for exchanges.
Relative to these efforts, Jasper’s main contribution is architecting for scale. Jasper adopts the idea of synchronized clocks from CloudEx, but scales the system much further by using a communication tree in both the inbound and outbound directions and augmenting this tree with in-network pacing and scheduling algorithms (§3). Additionally, Jasper also tackles the latency-security tradeoff on the “last hop:” the link from the exchange’s edge (e.g., gateway VM, proxy, SmartNIC) to the participant’s trading algorithm (§6). Jasper employs the widely used notion of outbound fairness (simultaneous delivery of data to MPs) from CloudEx.
3 Jasper Overview
Figure 3 presents an overview of Jasper. For communication between the exchange and participants, Jasper uses a bidirectional overlay tree: the exchange server is the root and MPs are the leaves, with intermediate proxy nodes. It is well known that trees help scale communication to many receivers [21, 22]. We build on such a tree to provide a multicast service for market data and handle bursty order submissions, but adapt the tree to the high-variance environment of the public cloud. Additionally, we adopt from CloudEx the idea of using synchronized clocks to (1) simultaneously disseminate market data and (2) sequence participants’ orders.
Outbound: ME to Participants. On the outbound side of an exchange, we augment the base tree with 3 techniques to lower multicast latency and minimize the latency variance: (i) round-robin packet spraying (RRPS), (ii) proxy hedging and (iii) receiver hedging. All 3 hedge the risk of some part of the system showing performance variance, as explained below.
To minimize the impact of latency spikes on the links between the tree nodes, proxies at each layer rotate their child VMs after each message, ensuring messages don’t always follow fixed paths. This approach limits the effect of latency spikes, as only a small subset of messages is impacted by any transient spike on any link within the tree.
We employ proxy hedging to mitigate the impact of straggler proxies that may introduce latency inflation. In this approach, each VM (a child proxy or a receiver) gets copies of a message from multiple proxies. Each VM processes the first received copy, discards the others, and forwards this copy to its children and some nieces. This way, if a proxy inflates latency for all messages going through it, its children can still receive market data promptly from an aunt proxy.
Lastly, if an MP’s VM becomes a straggler, it will not be able to compete effectively because of increased latency for its market data. We hedge this risk by allocating two VMs to each MP, where both run the same trading algorithm.
Inbound: Participants to ME. We use the same tree as above in the reverse direction for MP-to-exchange communication: MPs submit their orders to their parent proxies where it travels up the tree and reaches the ME at the root. Using a tree in the reverse direction reduces the fan-in factor of ME: the ME has to receive and process streams of orders from a small number of proxies instead of all MPs. This reduces packet losses as it reduces incast. We augment this reverse tree with in-network pacing and scheduling to efficiently handle bursty orders from MPs and relay them to the ME running at the root.
Packet pacing significantly reduces incast-style packet drops during periods of bursts. A proxy continuously senses One-Way Delay (OWD) between itself and its parent proxy. If OWD increases beyond a threshold, the proxy paces the packets to lower the OWD which in turn avoids packet drops.
However, pacing at proxies builds up queues, which are drained once a burst subsides. This can increase order latency and reduce ME throughput, which we compensate for via smarter scheduling. We develop Limit Order Queueing (LOQ) at each proxy, which prioritizes participant packets that contain critical orders, i.e., those closest to the mid-price. Critical orders are more likely to be matched upon reaching the ME, thus keeping the ME occupied without idling. Non-critical orders are likely to remain in the Limit Order Book (LOB) at the ME until the mid-price moves, and hence are deprioritized and wait longer at the proxies. LOQ increases the ME’s order matching rate and maintains low latency for matched orders during bursts.
Trust Models. Exchanges operate in an environment with strong incentives: participants have every incentive to act on market data as soon as it arrives and also have every incentive to set an earlier timestamp for their order. We develop multiple participant-to-exchange interfaces for the last hop between the participant’s trading algorithm and the leaf nodes of the tree. In the first interface, the leaf runs within a VM; the algorithm runs in a secure enclave within the VM. In the second, the leaf and algorithm run in separate VMs. In the third, the leaf and algorithm run in the same process.
These interfaces tradeoff performance vs. security: the first and second both don’t require the participant to reveal their algorithm to the exchange, but experience significant latency on the last hop. The third requires the participant to reveal their algorithm, at least in binary form, but has much lower last-hop latency. Between the first and second interfaces, we find that while their mean last hop latencies are comparable, the enclave-based solution has much lower variance.
4 Outbound Communication
On the exchange outbound side, market data is distributed to MPs fairly and with low latency. In on-prem exchanges, fairness—simultaneous data delivery to all MPs—is achieved by equalizing wire lengths between the exchange server and MPs and using low-jitter switches. Here, we develop overlay techniques atop a best-effort fabric to achieve fair and low-latency multicast that scales to a large number of receivers.
The simplest approach to transmit the same message to many receivers is multiple direct unicasts, where the sender sends the same message directly to each receiver. This increases the latency for later receivers: back-to-back unicasts are serialized over the sender’s egress, cumulatively increasing transmission and serialization delay.
Instead, we use an overlay tree to create a scalable multicast. Here, the sender sends messages to a few proxies, which relay them down the tree to more proxies, and recursively down to receiver VMs at the leaves. This proxy tree reduces the cumulative delay (Fig. 4) by limiting each node’s fanout. However, each additional tree hop introduces some latency. We minimize this by optimizing proxy packet processing with kernel bypass and zero-copy techniques.
Additionally, tuning the tree’s fan-out () and depth () is crucial for minimizing latency for a given number of receivers (). We conducted experiments with various <> for , and and found that setting and achieves near-minimum latency. Given the cloud’s inherent variability, more sophisticated tuning does not significantly improve latency, as detailed in Appendix A.
Additionally, we synchronize the clocks of all MPs, proxy nodes, and the ME using the Huygens clock synchronization algorithm [23]. As in CloudEx’s “hold-and-release” mechanism [4], we assign each multicast message a future timestamp called a deadline. After receiving a message, receivers hold it until its deadline before processing it to ensure fair dissemination of market data. The deadline is calculated to ensure all receivers likely receive a message before the deadline, and is set to the send time plus a conservative estimate of maximum latency from the sender to any of the receivers.
However, as the number of receivers increases and the latency shows high variance, the required holding duration by each receiver increases to ensure simultaneous delivery which inflates the multicast latency. When “hold-and-release” is used with our latency and latency variance reduction techniques, it scales gracefully as it only has to account for small residual latency variance thus requiring smaller holding duration to achieve fairness (§8.5).
We address the security issue of trusting the receivers to process messages at or after the deadline in §6. The implementation details of “hold-and-release” and the deadline calculation by scalably aggregating OWD estimates from all the receivers is presented in Appendix B. In the following, we describe techniques that enhance the proxy tree to reduce both latency and latency variance.
4.1 Round-Robin Packet Spraying
In round-robin packet spraying (RRPS), parent-child associations in the tree are flexible: proxies change their children after each multicast message. This provides two main benefits: (i) the effect of latency spikes is reduced and (ii) a sudden burst of market data is distributed among more network paths. Both benefits translate to reduced multicast latency.
In RRPS, for a given proxy, on every new multicast message, the set of children is circularly shifted by 1. For example, the first proxy in a layer will have the first proxies of the next layer as its children for message 0, the next proxies as children for message 1, etc. This ensures that each proxy has a parent proxy for each message, while the set of children for a parent continuously changes. With this technique, a single leaf node in a proxy tree has a total of different virtual paths starting from the root node. Successive messages traverse this abundance of virtual paths in a round-robin fashion. On the other hand, without RRPS, there is only one virtual path from the root to each leaf node and all messages traverse that path. The increased number of paths provides latency reduction benefits, as discussed below.
First, if a spike occurs on any VM-to-VM virtual link, only a subset of messages is impacted: because messages are round-robined across all available virtual paths, many messages will take alternative paths that avoid this virtual link.222Here, we assume that spikes are characteristics of individual virtual links, not VMs themselves—as is often the case. If spikes were a VM characteristic—as is also the case sometimes—many virtual links would simultaneously be affected, reducing the benefits of RRPS. For such straggler VMs, we use proxy hedging (§4.2).
Second, without RRPS, each VM communicates with a fixed set of children VMs for every multicast message. So a burst of messages has a higher chance of filling up queues in the network, inflating latency. However, RRPS uncovers and utilizes new virtual paths that were unused earlier. For example, consider two adjacent proxy layers: a parent layer with proxies and a children layer with proxies where each parent proxy has children. Without RRPS, the number of virtual paths utilized by messages going from parents to children is . With RRPS, there are virtual paths. This allows us to distribute bursts among many more virtual paths, which reduces queue build-up and latency.
At its core, RRPS works because there is an abundance of virtual paths available within an overlay proxy tree. This is because every VM in a cloud region can communicate with every other VM since a VM-to-VM link is virtual—effectively the network topology is a clique. This is unlike multicast trees in physical networks where the network topology is not a clique and is limited by physical constraints like geographic distance and cables.
4.2 Proxy Hedging
A VM’s performance in the public cloud can degrade due to factors like noisy neighbors or live VM migration, causing latency fluctuations for all messages passing through the VM. To mitigate this, we develop proxy hedging, where each node in the tree receives multiple copies of a message from proxies in the higher layer. The node processes only the first copy and discards duplicates, thus reducing the impact of any straggler proxy VM. This technique decreases both temporal and spatial latency variance, leading to more stable latency and reducing the delivery window size (the difference between the earliest and latest receiver’s latency). We first outline a basic design that offers latency improvements but results in high throughput wastage. Then, we introduce an optimized version that avoids excessive throughput wastage while maintaining latency benefits. For ease of explanation, we assume RRPS is not in effect, but the 2 techniques compose.
Strawman design: F+H Technique. Each (internal) tree node has extra children along with its original children, where is the hedging factor. For example, in a multicast tree with and (Figure 3), if we configure , we will add 1 hedge node in Layer-0 and 3 hedge nodes in Layer-1. Each hedge node will assist original proxies at the same layer and does not have children of its own. A hedge node assists these proxies by forwarding a message received from the hedge node’s parent to the hedge node’s nieces. In this way, a node receives copies of a message: 1 copy from its parent, and the other copies from the hedge nodes of the layer above it. This technique significantly reduces latency variance. However, it has a distinct drawback: Whereas a proxy only needs to multicast messages to the nodes, the hedge node needs to multicast the same message copies to nodes, creating an egress throughput bottleneck at the hedge node. We build on the observations to design the final scheme.
Final design: SiblingHedging. Instead of adding extra hedge nodes in a tree, we let each proxy node also take on the responsibility of hedging. Each proxy node sends messages to the children of of its siblings along with sending messages to children of its own. For example, in Figure 3 (left), when hedging is not enabled (i.e., ), proxy ( proxy in layer) only receives messages from . As a result, may suffer from high latency if or the path from to encounters latency fluctuations. By using hedging, not only receives messages from , but also receives the same messages from one (if ) other node, . With proxy hedging, each node has more than one parent. This technique shows significant benefits that we evaluate in §8.2.
With proxy hedging, a proxy processes the earliest received copy among copies of a message and discards the rest. This leads to significant latency reduction of each message, at the cost of some throughout wastage. We find out empirically as well as with Monte Carlo simulation that a small H (of up to 2) is enough for reaping most of the benefits of hedging. A Monte Carlo analysis of hedging is presented in appendix C that corroborates: (i) latency and its variance reduces as H increases, and (ii) increasing H shows diminishing returns.
Proxy hedging leads to redundant work which decreases the effective throughput (i.e., goodput) of a proxy. The goodput of a proxy node in SiblingHedging can be defined as: because each proxy sends redundant messages to the children of ’s siblings. One way to recover the lost throughput is to employ several parallel trees that share the root and the leaves (assuming the leaves have enough ingress bandwidth). The increased number of proxies also raise concerns of dollar cost, which indeed is a drawback here. The dollar cost can be ameliorated by utilizing spot VMs for proxies instead of on-demand VMs as they are siginificantly cheaper [24]. However, utilizing spot VMs requires proper fault tolerance that we address in §7.
4.3 Receiver Hedging
If a VM belonging to an MP becomes a straggler, the MP would not be able to compete with others as the market data it receives will lag behind other MPs. To hedge this, we provide each MP with two receiver VMs. Each receiver VM belonging to an MP runs the same trading algorithm, receives market data, and generates orders. If any one VM belonging to an MP suffers from performance variation or data is not disseminated fairly to it, then the MP’s other VM would help retain good performance, with high probability. This technique improves the performance guarantees significantly at the tail.
Receiver hedging doubles the number of receiver VMs, which increases the dollar cost. On the other hand, due to the scaling properties of the proxy tree, the increase in OML (overall multicast latency) is negligible when the number of receiver VMs doubles. Moreover, when we consider latency for an MP as the minimum latency experienced by any of its two receiver VMs, the OML improves (§8.4). Appendix F discusses how the two VMs of an MP stay synchronized.
4.4 Remarks on Multicast Packet Losses
In on-premises exchanges, the multicast service does not provide reliable delivery. Instead, a separate "rewinder" service is deployed that handles any retransmission requests from MPs [25, 26]. MPs use sequence numbers of the messages to detect gaps and request a retransmission. The exchange infrastructure is engineered to provide negligible packet losses so the MPs only have to contact the rewinder rarely.
The Jasper multicast service needs to similarly provide low losses. Our hedging techniques help lower the packet losses, but we observe that even without our hedging, the losses in outbound direction are small: we observe 0.007% lost packets in a benchmark on GCP while exhausting 80% of egress bandwidth (more details are in Appendix H).
5 Inbound Communication
The matching engine (ME) is located at the root VM of the tree, while orders come from leaf nodes (MPs/traders). Using the proxy tree in reverse for order submission helps reduce the submission load on the ME. Intermediate proxies gather orders from multiple downstream nodes and forward them upstream, so the root VM simultaneously processes packets from only proxies instead of MPs. The reduced fan-in decreases packet losses. In a benchmark with 10 VMs each submitting 5K orders per second to the ME, the packet loss rate is 1.58%. Adding two intermediate proxies to create a tree structure lowers this loss rate to 0.68%, as the proxies alleviate incast congestion by relaying orders to the ME.
A tree alleviates some incast congestion, but is not sufficient especially when order bursts occur. We use packet pacing at the proxies to tackle such bursts. Avoiding drops by pacing the packets comes at the cost of filling up queues at the proxies which may increase order submission latency. To maintain low latency for orders and achieve high order matching rate of ME, we design a distributed packet scheduling policy, implemented at the proxies. The policy, realized through a priority queue, Limit Order Queue (LOQ), prioritizes orders that have prices closest to the current mid-price.
5.1 Pacing Outgoing Packets
Proxies continuously request the OWD estimates of packets sent to parent proxies or the ME. An increasing OWD is an indicator of congestion at the receiver (i.e., parent proxy or ME) and the proxies use this to slow down – increase the interpacket duration – to avoid packet drops.
Aggressive pacing can lead to bandwidth underutilization, while insufficient pacing risks excessive packet drops due to congestion. Ideal pacing balances these concerns by optimizing bandwidth usage, and adjusting pacing to avoid drops during bursts while reclaiming bandwidth once bursts vanish. In Jasper, prioritizing drop prevention is more important than operating at the maximum bandwidth. Sharding techniques for MEs can improve throughput [4], hence our focus is on minimizing packet drops during bursts. Nonetheless, it is favorable to utilize as much of the throughput as possible if it does not introduce substantial packet loss.
Our pacing heuristic calculates an amount of time to wait before sending the next packet (the pacing quantum) based on how much the most recent OWD estimate () deviates from a threshold OWD . is computed as the median OWD of all the packets sent in the last . The heuristic increases the pacing quantum by if increases by over . And the quantum is decreased by if falls below by . The heuristic is aggressive in increasing the pacing as compared to when decreasing the pacing because it is critical to avoid packet drops. When is sensed to be decreasing below , we slowly (decrease by vs increase by ) decrease the pacing to reclaim the available bandwidth. A sudden decrease is avoided to account for imperfect estimates. After pacing is applied, we empirically observe that we maintain a stable throughput (orders received per second) at the ME, despite being confronted by bursts of orders. We also observe that pacing significantly reduces packet drops. In §8.7, we show pacing can achieve a reasonable balance between throughput and packet losses. Our attempts to further improve this simple pacing heuristic were not fruitful because of the noise inherent in calculating as a cloud tenant. It is possible that there are better pacing schemes that Pareto dominate our scheme on throughput and loss; for now, though, we have found our simple scheme to suffice.
Parameter Selection We pick and the increase and decrease amounts offline by running a control experiment where VMs communicate at a fixed rate which is expected to be the equilibrium rate for order submission.
5.2 Limit Order Queue (LOQ)
LOQ implements a distributed scheduling policy across all proxies of the overlay tree, improving ME’s order matching rate and reducing order latency. Its benefits are pronounced when packet queues form at the proxies.
LOQ leverages the observation that orders with prices far from the mid-price tend to wait in the LOB, even if they arrive earlier than orders closer to the mid-price. These noncritical orders can afford to remain in proxy queues without impacting the matching process. If a proxy knows the mid-price, it can prioritize outgoing orders based on their proximity to this price. Critical orders—those closer to the mid-price and more likely to be matched—are relayed to the ME first. This prioritization boosts ME throughput (orders matched per second) and reduces the latency of matched orders.
The above approach to LOQ relies on proxies having access to mid-price information. One way to provide this information is for the ME to periodically multicast the mid-price to all proxies, which would increase the ME’s workload, especially during times of high activity and order overloads. Another solution is for each proxy to maintain an approximate LOB based on the orders it relays to the ME. However, this adds the overhead of maintaining an LOB at the proxies and may diminish the benefits of LOQ if a proxy’s LOB significantly diverges from the ME’s LOB.
Mid-price agnostic LOQ: To avoid maintaining mid-price information at each proxy, we design LOQ to be agnostic of the mid-price information. A proxy needs to prioritize some orders over others based on their probability of getting matched as they arrive at ME. The distance to the mid-price is a good indicator of this probability. However, a proxy does not need the exact probability but just the ordering relationship i.e., whether the probability of one order is higher or lower than some other order that it has in the queue. We leverage this to design a mid-price agnostic LOQ as follows.
LOQ at each proxy consists of two separate priority queues, one for bids and one for asks. The bids queue sorts the orders by price in descending order while the asks queue sorts the orders by price in ascending order. This reflects the ME behavior that the highest bid and lowest ask are selected for a potential match. A proxy consults the two priority queues in a round-robin fashion for dequeuing the orders and forwarding them to ME. This design eliminates the need for mid-price information at the proxies, simplifying the system.
LOQ also needs to respect price-time-priority that an ME uses for order matching. Each priority queue at the proxies first sorts the orders by price, as discussed above, and then by order generation timestamp, . This ensures that when a proxy dequeues an order from LOQ for relaying to ME, another order with the same price and lower does not exist in the queue. A sequencer delay is used by ME to account for any network reordering: An ME only dequeues an order for processing if it was received at least ago which gives enough time, with high probability, to the other orders to arrive and get prioritized if they deserve to do so e.g., they have a lower or an LOQ deemed it to arrive at ME earlier.
5.3 Deconstructing LOQ
LOQ improves order matching rate and reduces latency of matched orders by reordering queues at the proxies. A natural question is whether we are sacrificing correctness by such reordering. The reordering only aims to prioritize orders with a higher likelihood of getting matched, i.e., those with a lower chance of remaining in the LOB and idling the ME upon arrival at the ME. Intuitively, it seems LOQ does not affect the matching behavior of ME, but we have not formally shown that. Empirically, we use inbound fairness333Inbound fairness is distinct from the (outbound) multicast fairness to quantify how much LOQ changes the behavior of ME.
Inbound fairness: For an ME, it’s critical to ensure that orders generated earlier are matched before those generated later if both provide the same market value. This concept of inbound fairness is a key feature of matching algorithms like price-time priority. Any infrastructure for exchanges needs to ensure an ME can do the above: if an order submitted earlier () should be matched before a later order () then the ME should have received both by the time it chooses for execution.
CloudEx [4] proposed inbound unfairness ratio to quantify the inbound fairness. The ratio is calculated as the percentage of sequenced444a sequencer sequences all the orders based on their generation timestamp orders received out-of-order by the ME. We find that LOQ shows very poor inbound unfairness ratio (, where ideal value is 0%). Upon interrogation, we realize the CloudEx’s definition is unnecessarily strict. It may increase unfairness ratio under a benign scenario: an order with order generation timestamp, arrives later than with where but has a price much closer to the mid-price while has a price farther away and would only wait in the LOB even if it arrives at ME earlier than . This is the exact scenario which LOQ takes advantage of.
We realize the issue with CloudEx’s definition, and propose a new one: order matching fairness. Assume MPs attach with each order they submit.555Security issues with timestamping are addressed in §6. In order matching fairness, an order with and price is only considered unfairly treated if it arrives at the ME and an order with and price has already been matched by ME where and . The percentage of unfairly treated orders among the total matched orders constitute inbound unfairness ratio.
By focusing on the matching of orders and price, the order matching fairness allows packet scheduling policies to send orders, to ME, with different prices out of sequence to enhance metrics like the order matching rate. We believe our new inbound fairness definition captures the true intent of inbound fairness, however we lack a formal analysis, which is left for future work. In §8.8, we find that Jasper leads to small inbound unfairness ratio which remains unchanged with or without LOQ, suggesting that LOQ may not affect correctness.
6 Jasper Trust Models
In an ideal trust model, an exchange wouldn’t need to rely on MPs to adhere to protocols like accurately time-stamping trade orders or withholding market data processing until a set deadline to ensure fairness. Likewise, MPs wouldn’t have to disclose their trading algorithms to the exchange. These guarantees should be achieved without incurring any performance overhead, such as increased latency or reduced throughput.
An ideal model should also not introduce any jitter for packets going from the hold-and-release (the exchange’s program to delay release of market data) to MPs’ trading algorithms. Fairness is achieved at the level of the hold-and-release program using CloudEx’s time synchronization mechanisms, which Jasper also leverages. To ensure fairness at the level of trading algorithms, there should be a constant latency between the exchange program and the trading algorithm. Achieving this ideal model is impractical due to the tension between security and performance that we present. We propose three trust models (Fig. 5), discuss their respective trade-offs, and offer our recommendation.
Model 1: MPs give their programs to the exchange. Exchange controls the VMs where the MPs’ trading programs run. The exchange runs hold-and-release in these VMs and then forwards the multicast messages to the MPs’ programs. No significant latency or throughput overhead is incurred. The jitter between the exchange’s program and the MP’s program can be minimized to be non-significant () as forwarding messages between the two programs can be mere function calls. However, MPs have to reveal their proprietary trading programs to the exchange in this model.
Model 2: Separate gateways for hold-and-release. Model 2 deploys gateways between the exchange’s infrastructure (ME and proxy tree) and the VMs where MPs’ algorithms run. The exchange controls the gateways hosting the hold-and-release programs, while MPs control the VMs running their algorithms. It avoids the need for MPs to reveal their programs or for the exchange to trust MPs with hold-and-release. However, it introduces latency between the hold-and-release program and MPs’ programs, as they run in separate VMs. Throughput is unaffected, depending on VM bandwidth, but OWD between VMs is around , with high jitter due to cloud latency fluctuations and spikes. In the absence of the spikes, jitter may be in the tens of microseconds. Even with simultaneous delivery at the gateway level, fairness at the receiver VM level is not guaranteed due to this jitter.
Model 3: Trading programs run in secure enclaves Model 3 leverages the confidential computing capabilities of VMs equipped with secure enclaves such as AWS Nitro Enclave. An enclave can only talk to its associated (parent) VM (and AWS Nitro Hypervisor). The exchange owns the VMs,666Appendix D shows a trivial but under-performing alternative model, subjecting incoming messages to hold-and-release before forwarding them to the respective enclaves within the VMs. The MPs’ program executes within the enclave.777Loaded via remote attestation mediated by AWS Nitro Hypervisor MPs do not have to reveal their trading programs to the exchange and the exchange does not have to rely on the MPs to run hold-and-release. Although perfect security boundaries are achieved, performance overhead is incurred: high latency, extremely low throughput, and non-negligible jitter.
The latency between a VM and an enclave is similar to the latency between two VMs, making it comparable to Model 2. However, Model 3 benefits from reduced jitter between a VM and its enclave, likely because communication between them does not traverse the network. Optimizations described next further reduce the jitter. Nonetheless, throughput between a parent VM and its enclave is significantly lower (about 70% lower) compared to the parent VM’s ingress, as noted in [27].
Reducing jitter between a VM and an enclave We reduce the latency variance between a VM and its associated enclave with some optimizations: isolating CPUs, reducing scheduling-clock ticks, and pinning threads to cores. Figure 6 shows the 90th percentile latency between a parent VM and its associated enclave for each tumbling window of . After the optimizations, the latency becomes much more predictable. We observe a jitter (the difference between and latency) of .
Recommended Model. Optimal performance is attained with Model 1, though MPs need to reveal their programs to the exchange. If an exchange is obliged to not reveal its clients information to third parties (potentially binding due to Section 6801 of 15 U.S. Code in some territories [28]), this model should be adopted in practice. We use Model 1.
7 Fault Tolerance
Domain | Fault Type | |
Transient | Long Term | |
VM | {Proxy, Receiver} Hedging | {Proxy, Receiver} Hedging + FD |
Link | Proxy Hedging + RRPS | - |
Table 1 summarizes how different types of faults are guarded against in Jasper. We do not expect long-term VM-to-VM link failures in the cloud. Similarly long term failures of on-demand VMs are not expected [29]. However, failures of spot VMs need to be handled if we use low-cost spot VMs to decrease the dollar cost of proxies [24]. However, if all proxies are spot VMs, then the entire operation of an exchange may cease as spot VMs exhibit all-or-nothing availability [24]: if one spot VM becomes unavailable then it is expected that several other spot VMs will become unavailable. We may need to boot and utilize on-demand VMs until spot VMs becomes available again. Time to boot VMs vary depending on the virtualization stack and it is common to have boot times between 1s and 60s. However, starting a stopped VM is quicker than booting a new VM [30]. Empirically, we find that starting 100 stopped c2d-highcpu-8 VMs on GCP takes . So an exchange may need to stop its operation for several seconds upon any spot VMs failure.
An alternative model designates every even-numbered proxy at each tree level as a spot VM, with the others as on-demand VMs. This setup offers cost benefits and enhances performance during failures. Proxy hedging ensures that each proxy communicates with the children of another sibling (if ). If all spot VMs become unavailable, the messages still reach their destinations because neighboring odd-numbered proxies will be active and handling communication through proxy hedging. Although this may introduce latency fluctuations, it is preferable to an approach that completely disconnects the exchange and MPs upon failures. Additionally, this model supports an efficient failure detection (FD) mechanism, as each on-demand VM can regularly check its sibling spot VM, with checks occurring frequently e.g., every .
8 Evaluation
We answer the following questions.
How does multicast latency compare with AWS TGW-based multicast? §8.1
- (2)
How much outbound fairness is achieved? §8.5
How does latency and fairness compare to CloudEx? §8.5
How scalable is Jasper? §8.6
How effective is packet pacing in minimizing packet drops during bursts? §8.7
What impact does parameter have on packet drops and the achieved throughput? §8.7
How much does LOQ help in enhancing ME’s throughput and lowering orders’ latency? §8.8
Does LOQ affect order matching fairness? §8.8
What are the trade-offs in the Jasper trust models? §8.9
Unless mentioned otherwise, we use 100 MPs and use 5K multicast messages per second (MPS). Each MP has a separate VM (or two VMs in an experiment for receiver hedging). The overall multicast latency (OML) refers to the latency experienced by the last receiver that receives a multicast message. The delivery window size (DWS) is the maximum difference in the latency of any two receivers. We use the Huygens algorithm [23] that synchronizes clocks of all the VMs with a offset of .
8.1 Multicast Latency And Its Comparison
We consider the following baselines and compare their multicast performance with Jasper.
Direct Unicast (DU): When conducting multicast, the sender directly sends a copy of a multicast message to each receiver. We use io-uring to create a strong baseline as we observe it reduces the latency by minimizing the overheads of syscalls when doing batch I/0.
AWS Transit Gateway: AWS-TGW-based multicast is provided provided by AWS [1]. It requires the sender to send the message to a gateway which then replicates and sends one copy to each receiver. It can support at most 100 receivers.
Figure 8 shows that Jasper distinctly outperforms DU and AWS-TG. This experiment was performed entirely on AWS. The median latency for Jasper is which is lower than the latency of AWS TG () and lower than the latency by DU (). Moreover, Jasper shows predictable latency as higher percentiles are very close to the median in contrast to the other techniques.
8.2 Proxy Hedging
Reduced overall multicast latency. Figure 8(a) compares the OML CDFs under different hedging factors (). represents the case when proxy hedging is not enabled. Compared with , yields distinct latency reduction as the CDF curve is shifted to the left. The latency reduction is marginal as H increases from 1 to 2. We show later that setting yields better DWS than .
Reduced temporal and spatial latency variance. We calculate the median OML over a tumbling window of 5000 messages to study the temporal latency variance during the multicast. Figure 8(b) shows Jasper with hedging exhibits lower temporal latency variance compared with the non-hedging scheme (). Enhancing the temporal predictability of latency, even for short time-steps, is beneficial in multiple ways, particularly in the hold-and-release mechanism, which relies on latency estimates to determine message deadlines.
Figure 8(c) compares the CDFs of DWS for Jasper with different hedging factors. Hedging substantially reduces the spatial latency variance (i.e., it shrinks the DWS of multicast messages). The 99th percentile delivery window size is with but with . However, as grows from 1 to 2, the window size reduction, although it still exists, becomes less distinct.
Reduced OWDs for each receiver. Proxy hedging improves OML as the OWD to each receiver is reduced. Figure 10 shows OWD for each receiver with () and without hedging. For this experiment, we interleave hedged and non-hedged messages to precisely study the impact on OWD. With interleaving, any latency fluctuations on the VMs or the paths impact both types of messages. All even-numbered messages are not hedged while odd-numbered messages are hedged. After the experiment, we separate messages based on their parity and plot Fig. 10: receivers’ OWD improves after hedging.
Cost of Proxy Hedging. Proxy hedging lowers multicast throughput in proxy VMs due to redundant tasks like sending messages to nieces. Reclaiming this lost throughput requires parallel proxy trees with a shared root (sender) and leaves (receivers), assuming receivers can handle the traffic. Table 2 details the cost of proxies on AWS, based on the c5.2xlarge instance at $0.34/hour for 100 and 1000 multicast receivers which is minuscule compared to on-premises exchange’s infrastructure cost and colocation fees [31, 7].
N | # of proxy VMs | # of proxy VMs | Cost/hour |
(H=0) | (H=2) | (H=2) | |
100 | 10 | 30 | $10.2 |
1000 | 110 | 330 | $112.2 |
8.3 Round Robin Packet Spraying
We evaluate the impact of RRPS on the multicast latency. We utilize a base multicast message rate of 10K MPS, 125 receivers and a burst of occur every . During the burst, the message rate becomes . Table 3 summarizes the results: multicast latency decreases when RRPS is used. On GCP, we see an OML reduction of , however, on AWS the reduction can approach . We also note that with low message rates, the OML reduction is not significant: with message rates MPS. This happens because the inter-packet duration is small enough that any transient latency spike is not able to impact consecutive messages so RRPS does not help much by re-routing messages.
Setup | On AWS | On GCP | ||
50p | 90p | 50p | 90p | |
Improvement due to RRPS (%) | 15.12% | 69.35% | 9.83% | 9.18% |
RRPS show improvements in OML when message bursts are introduced as it can distribute the bursts among several network paths. Without bursts, the OML reduction still happens but is less than on both AWS and GCP. Additionally, RRPS can help reduce OML deterioration caused by latency spikes on paths between VMs, as no single path is used for consecutive messages. Evaluating this aspect is difficult as we cannot inject spikes on the paths and is left for future work. A simulation-based result for this aspect would trivially demonstrate the desired outcome and is omitted.
8.4 Receiver Hedging
The receiver hedging improves OML (Table 5) and achieves fairness (i.e., DWS at , when hold-and-release is used). Receiver hedging is not utilized in any experiments other than Table 5. This technique improves Jasper performance at the cost of one extra VM per MP. Doubling the amount of receiver VMs () sometimes also require restructuring the proxy tree which incurs extra dollar cost: 10 proxies are required for =100, 14 proxies for =200.
RH | OML () | ||
50 | 99 | ||
100 | No | 139 | 248 |
200 | No | 141 | 243 |
100 | Yes | 99 | 146 |
T | Loss | Thpt. | |
Ideal | - | 0% | 80K |
No Pacing | - | 3.1% | 78K |
Pacing | 80 | 1.6% | 76K |
Pacing | 75 | 0.7% | 62K |
Pacing | 70 | 0.3% | 34K |
N | P(F) (%) | OML |
100 | 92 | 175 |
1000 | 84 | 238 |
8.5 Simultaneous Delivery
Fairness of data delivery among the market participants (MPs) is crucial for HFTs. We adopt the same fairness definition as used by on-prem financial exchanges: every MP should receive a multicast message almost simultaneously i.e., DWS should be negligible. In on-prem exchanges, the DWS is measured in nanoseconds, however, in the public cloud, the problem has not been solved even at the level of microseconds. Jasper, thus tries to achieve a DWS of . We compare with CloudEx [4] which uses the hold-and-release mechanism but does not use a proxy tree or any overlay techniques for improving multicast performance.
Figure 11(a) shows that Jasper can achieve a DWS of at very high percentiles (up to ). Without proxy hedging, the DWS becomes larger at an earlier percentile. CloudEx achieves fairness but at the cost of high OML.
Figure 11(b) shows that OML increases significantly for CloudEx and Jasper with no proxy hedging. In these systems, the deadlines calculated by the hold-and-release are far into the future to cover the high latency variance. This leads to high holding duration at the receivers for CloudEx and Jasper without hedging as shown in figure 11(c), increasing OML.
8.6 Scaling Jasper Multicast
Scaling N to 1000. In contrast to the previous fair multicast solutions that only work with a few 10s of receivers (e.g., [7, 5, 1]), Jasper aims to implement a more scalable multicast service which can support 1000 receivers and potentially even more. Figure 10 plots Jasper’s multicast latency for 100 receivers and 1000 receivers. Our overlay techniques for reducing latency enable Jasper to keep a graceful latency growth as the number of receivers increase.
Fairness for an increasing number of receivers. When employing hold-and-release, a median DWS of is achieved for N = 1000. However, the highest percentile at which DWS stays decreases as increases. We define the probability of fairness (P(F)) as the highest percentile where DWS is . Table 6 shows OML and P(F) for N = 100 and 1000.
Throughput. With our current implementation, we can exhaust 80% of the egress bandwidth of a c2d-highcpu-8 VM on GCP while maintaining negligible packet losses. The specified VM has an egress bandwidth of 16Gbps. With 466B packets, we achieve a multicast message rate of 350K MPS, without proxy hedging. The OML stays the same as we increase the message rate from 5K to 350K, but increases . In the experiment, we have 100 receivers and 10 proxies, with no proxy hedging. The sender and each proxy replicate each message times hence, with 350K messages per second, the sender generates 3.5M packets per second, utilizing 81.5% of the available egress. The proxy hedging would reduce the multicast rate according to the factor as each proxy replicates each multicast message times e.g., with , we achieve 175K messages per second. For an optimized implementation, we use DPDK and Zero Copy packet processing explained in Appendix E.
8.7 Packet Pacing At Proxies
Pacing packets at the proxies in response to increasing OWD helps minimize the packet losses. Our technique for pacing tries to achieve good throughput while minimizing packet drops. We evaluate it in an experiment with 10 MPs, 2 intermediate proxies, and a ME VM organized in a tree structure. Each MP submits 7K orders per second to its parent proxy, which relays it to the ME. We introduce a burst of orders after every where each MP’s order submission rate doubles. Table 5 shows pacing decreases loss rate (# of pkts lost /s), but incurs throughput (# of pkts received by ME /s) loss.
The effectiveness of pacing is influenced by the parameter , which estimates the one-way delay OWD between two VMs. If is set too low, pacing increases unnecessarily, causing a decrease in throughput (packets received by ME per second). Conversely, if is set too high, the pacing fails to increase adequately in response to rising OWD, resulting in significant packet losses.
Jasper currently uses UDP for order submission, which limits our ability to compare with a scenario where MPs employ TCP congestion control and retransmit lost packets. We plan to address this limitation in the future. One immediate drawback of having MPs handle retransmissions is that packet queues would form at the MPs rather than at intermediate proxies. These queues at MPs are less likely to benefit from LOQ, as the orders in the queue may lack price diversity.
8.8 Limit Order Queue (LOQ) Performance
We compare LOQ to a simple priority queue, which we refer to as SimplePQ. Unlike LOQ, SimplePQ does not differentiate between critical and non-critical orders; it merely prioritizes earlier orders over later ones, similar to how LOQ handles orders with the same price. In this experiment, we have 10 market participants (MPs), each submitting 7K orders per second. Two intermediate proxies, running LOQ or SimplePQ, relay orders from the MPs to the ME. Order bursts occur, with each MP doubling their order submission rate every . Bursts lead to queue build ups at the proxies (because of pacing), where LOQ/SimplePQ operate. As shown in Figure 13, LOQ outperforms SimplePQ in ME’s order matching rate and matched orders’ latency. Thin lines represent the orders received per second by ME (pacing is in effect) which are similar (but noisy) for both SimplePQ and LOQ. Thick lines represent order matching rate and orders latency in their respective plots showing LOQ’s better performance.
The advantages of LOQ stem from the presence of noncritical orders, which are kept in the proxy queues longer, allowing the available egress to be used for critical orders. We examine the benefits of LOQ under different ratios of critical to noncritical orders. In this experiment, MPs uniformly sample bid/ask prices from a predefined range to generate orders. We adjust the portion of the bid price range that overlaps with the ask price range. These overlapping segments result in orders that get matched at ME, while other orders remain in the LOB. Figure 13 shows as the ratio of overlapping price levels to the total number of levels decreases, the benefit of LOQ increases. We plot the metrics for the bursts duration which are calculated from the orders matched during the bursts that we inject. A ratio of indicates that there are total price levels from which MPs sample prices to generate orders, but only price levels lead to immediate matches at ME. As increases while remains constant, the number of critical orders decreases. LOQ identifies these critical orders and prioritizes them for processing by the ME before non-critical orders. Consequently, LOQ demonstrates a 150% increase in ME throughput and a 99.9% reduction in latency for matched orders when888 is commonly used in the literature [32] . To conclude this section, we repeat one experiment and increase MPs from 10 to 1K with and observe that LOQ shows benefits over SimplePQ: 98% decrease in the latency and 10% (86%) increase in overall (during bursts) matching rate.
How much order matching fairness is achieved? In the previous experiments with MPs, we achieve an unfairness of less than 5%. For the experiment with 1K MPs, we achieve an unfairness of 13%. The ratio stays similar with LOQ and SimplePQ . We only attempt to show here that LOQ does not alter the fairness ratio and the ratios are small enough that a substantial majority of orders are treated fairly. The absolute value of fairness depends on sequencer delay that presents a trade-off between latency of orders and fairness. More the sequencer delay, more ME waits before processing an order to counter any reordering by network, thus higher the latency. This aspect is orthogonal to Jasper and is studied in CloudEx.
Ideal | Model 1 | Model 2 | Model 3 | |
Latency () | ||||
Jitter () | ||||
Through- | No | No | No | |
put | Change | Change | Change | Reduced |
8.9 Performance and Security Tradeoffs
Jasper trust models are evaluated based on latency, jitter and, throughput overheads. We send 8K messages per second from the hold-and-release programs to the MP’s programs for 30 minutes. Jitter is calculated by measuring the difference between and . A large jitter shows that even if simultaneous delivery/fairness is achieved at the level of hold-and-release programs, the messages experience jitter before reaching MPs which may ultimately affect the fair delivery.
In model 2 and 3, the exchange and the MPs have a strict security boundary where (i) the exchange does not rely on MPs for the correctness of its operations and, (ii) the MPs do not have to reveal their trading algorithms to the exchange. In model 1, MPs are required to handover their proprietary trading algorithms to the exchnage so that it may run them and correctly follow its protocols i.e., hold-and-release. In model 2, one could assume that the the last layer of proxies serve as the gateways so no extra latency is added. However, the hold-and-release and trading programs would still be away and susceptible to latency fluctuations.
Table 7 summarizes the results of evaluating all 3 models. Model 1 outperforms others in performance. Model 3 is better than model 2 in jitter while model 2 is better in throughput. Model 1 already achieves fairness at the level of trading programs while model 3 can be made to do so by artificially inflating the hold-and-release deadlines by to counter the jitter. For fairness with model 2, a large inflation would be required (due to latency spikes) and can be accurately found by an offline trial and error method.
9 Related Work
Multicast: Prior works on application level multicast [33, 22, 34, 35, 21] mainly focus on finding optimal paths in a network (or overlay mesh) using cost models for network links that capture heterogeneous link bandwidth or latency characteristics. By contrast, Jasper focuses on (i) minimizing the delays incurred at hosts while transmitting multiple message copies, and (ii) achieving a small latency difference between the receivers. On the other hand, switch-based multicast [36, 37] is not available to tenants of the public cloud due to scalability issues [38]. Some clouds provide special offerings for multicast [1] that is indeed implemented using multiple unicasts and lack performance. We compare against and outperform AWS TGW-based multicast in terms of latency and scale.
Collective Communication: Collective communication [39, 40, 41] employs overlay trees for broadcast and all-reduce [42], but typically supports fewer than 100 receivers. For example, Hoplite [39] optimizes bandwidth for nodes using a tree-based topology. Jasper focuses on minimizing outbound latency and variance, handling bursts, and maintaining low order latency for larger scales, i.e., 1K nodes.
Pacing for Transience: OnRamp [43] introduced the idea of decoupling transience and equilibrium management in congestion control. Pacing was introduced to counter transient congestion. While similar in nature, Jasper’s pacing happens the intermediate proxies in a hop-by-hop fashion, instead of at the end-hosts in end-to-end fashion like in OnRamp.
10 Conclusion
This paper introduces Jasper, a scalable financial exchange hosted in the cloud. Jasper features a low-latency and fair multicast service, as well as an order submission service designed to gracefully manage order bursts. We open questions that warrant more interrogation: (i) what is a necessary and sufficient definition for inbound fairness, (ii) does order matching fairness respect continuous price-time priority algorithms’ semantics? This work does not raise any ethical issues.
Appendix A Appendix: Deciding D and F for multicast tree
To understand how the end-to-end latency varies as and grow, we conduct a series of experiments with various numbers of receivers (). Table 8 displays the latencies for different configurations of <> for a given number () of receivers. We see that the latency reaches a minimum point at different depths for different values of . At a small scale (), increasing does not bring latency benefits. As the scale grows from to , the benefits of increasing while reducing grow because the reduced message replication delay and transmission delay outweigh the added overhead of new hops in the path of messages. However, as goes beyond a certain threshold (e.g., grows larger than 3 when ), the latency is not reduced but gets worse. Based on our experiments, we find fixing usually leads to a desirable to generate a multicast tree with low latency. We establish our heuristic rule to construct the multicast tree as follows: Given receivers, we fix and then derive (round to the nearest integer). We also tested a more sophisticated alternative mechanism as described in [15], and did not observe a significant performance boost compared to our heuristic.
In our testing, we find that variance in VM performance in the cloud [44] makes it challenging to differentiate the “optimal” values of and from values picked using the aforementioned heuristic. We observe through repeated experimentation that minor changes in and do not show a significant performance difference with high statistical confidence. So it is sufficient to select and in the neighborhood of the optimal value, which is why our heuristic is effective. We further find that a unit increment in comes at the added latency of and a unit increment in adds per layer.999Using our high-performance implementation on a c5.2xlarge VM in AWS. Message size is 460B. A tree constructed using linear models based on these unit increments performs comparably to the tree constructed using our heuristic of maintaining close to 10 and . An example scenario for deciding tree structure: For N=100, F=10, D would be 2. For N=200, F=10, D would still be 2 and F would need to adjusted accordingly to support all 200 receivers. So F would come around to be be 14 as .
Appendix B Appendix: Scalable Simultaneous Delivery
Financial trading needs to ensure the fairness of the competition. Fairness in data delivery [4, 5] means that the market data from the exchange server should be delivered to every MP at the same time so that an MP may not gain an advantage over the others during the competition. While there are also some recent works trying to alter the definition of fairness [7, 20], these variant definitions require more research before they are confidently adopted by the exchanges. Therefore, Jasper uses the original definition of fairness employed by on-premises financial exchanges. In short, in Jasper the fair delivery of data means simultaneous delivery of market data to all the multicast receivers.
Realizing perfect simultaneous data delivery to multiple receivers is theoretically unattainable [45]. Nevertheless, Jasper tries to empirically minimize the spatial (i.e., across receivers) variance of the latency of messages. Our hedging design (§4.2) has created favorable conditions to minimize the spatial variance. By using hedging, Jasper can achieve consistently low variance for each receiver over time, and the spatial latency variance across receivers is kept low. Beyond this, we employ a hold-and-release mechanism (by using synchronized clocks) to eliminate the residual spatial variance at the end hosts and enforce simultaneous delivery across receivers. The hold-and-release mechanism was introduced by CloudEx [4] but it does not scale well and leads to high end-to-end latency (§8.5). We describe a modified hold-and-release mechanism that helps us scale further, while maintaining a low end-to-end multicast latency.
Hold & Release mechanism. To implement the hold-and-release mechanism, Jasper leverages the accurate clock synchronization algorithm, Huygens [23], to synchronize the clocks among the sender and receivers. Receivers keep track of the one-way delay (OWD) of the messages received from the sender. Each receiver takes the 95th percentile of its OWD records at regular intervals and sends the results back to the sender using an all-reduce mechanism explained later. The sender calculates the maximum of these OWDs called Global OWD for messages using the gathered statistics: where is the OWD estimate reported by the -th multicast receiver.
Once an has been calculated by the multicast sender, it attaches deadlines to all outgoing messages. A deadline is calculated by adding to the current timestamp when sending a message. Upon receiving a message, a receiver does not process it until the current time is equal to (or exceeds) the message’s deadline. This mechanism leads to almost simultaneous delivery, modulo clock sync error. In §6, we discussed possible security mechanisms to ensure that a receiver (an MP) waits until its deadline to process a message, when it’s in the MP’s self interest not to wait.
Deadlines all-reduce. In the design of Jasper’s hold-and-release mechanism, all the receivers have to send estimates of the OWD they experience back to the sender so that the sender can estimate the deadlines for the subsequent messages to multicast. If all the receivers send their estimated OWD directly to the sender, incast congestion occurs at the sender, leading to increased message drops.101010We measure that more than 20% of the packets are dropped when 100 receivers attempt to send OWDs back to the sender simultaneously. To avoid the high volume of incast traffic, we reuse the multicast tree to aggregate the OWD estimates in an all-reduce manner [42]. Specifically, each receiver periodically sends its OWD estimate to its parent proxy. As each proxy has a limited number of children, we do not risk in-cast congestion here. Each proxy (i) continuously receives estimates from its children; (ii) periodically takes the maximum of all the received estimates, ignoring some children OWDs if they have not yet sent in an estimate, and; (iii) sends the max value to its parent proxy. In this way, the sender at the root calculates deadlines for messages by only receiving the aggregated estimates from the first proxy layer instead of all the receivers.
Appendix C Appendix: Hedging Analysis Via Monte Carlo
When VM hedging is enabled, we may model the latency experienced by a receiver as follows.
We use function to represent the latency from node to node . We use to represent the root node (i.e., the sender) in Figure 3, and to represent the node (i.e, a proxy or a receiver) in Layer . Then, given a node , the end-to-end latency from the root sender to this node can be recursively defined as the following random variable ( denotes uniform random distribution):
Each is assumed to be IID. However, the latency is impacted by the order in which a parent proxy sends out the messages. This happens because of the replication and the transmission delay of sending messages. If the number of downstream nodes for a proxy stays small, we can ignore this order and assume to be IID.
Achieving a low variance of would mean that the latency over time may not deviate much from the expected value, helping in achieving consistent latency over time. It also shows that different VMs (at the same level of the tree) in Jasper may not experience latency significantly different from each other, reducing the difference between the maximum and minimum latency experienced among all the receivers for a multicast message.
We run a Monte Carlo simulation of random variable with different values of and to understand its behavior. The simulation is run for 100k iterations. Based on the simulation results (Figure 14), we have three main takeaways.
No Hedging High Latency Variance: With no hedging, the depth of a tree and latency variance are directly correlated. Figure 14(a) plots the CDF for and shows mean value and standard deviation for different configurations of the multicast tree. As grows larger, we see both and increase distinctly, indicating that just a proxy tree (i.e., no hedging) suffers from more latency variance when the tree scales up.
Hedging Low Latency Variance: Our hedging makes the correlation between and become less significant. In Figure 14(b), we can see the latency distribution becomes narrow (i.e., reduced ) as grows from 0 to higher values.
Low Overall Latency & Diminishing Returns on : In Figure 14(c), we fix and keep increasing . Figure 14(c) shows that hedging not only reduces the latency variance (i.e., the distribution becomes narrower), but it also helps to reduce the overall latency (i.e., the distribution moves leftwards). However, as grows larger, the incremented performance gains diminish, and most performance improvement is obtained when grows from 0 to 1. It shows that a small value of (>0) is enough to reap the benefits of VM hedging. This is useful because a small saves bandwidth.
Appendix D Appendix: Trust Model 4: Hold-and-release runs in a secure enclave
An alternative trust model is that a receiver VM is governed by a trader while a hold-and-release program runs in the secure enclave. However, to ensure that messages are only processed by traders after they go through the hold-and-release program, the exchange needs to encrypt the messages. A receiver VM receives such messages, forwards them to the enclave, and then receives decrypted messages at a later time as defined by the hold-and-release program. This method of utilizing enclaves adds encryption/decryption overhead without providing any benefits over Trust Model 3, so this is only presented for completeness.
Appendix E Appendix: Optimizations for High Throughput
Decoupling DPDK Tx/Rx processing: We leverage the high-performance lockless queue [46] provided by DPDK to decouple the Tx/Rx processing logic. On each 8-core proxy VM in Jasper, we allocate one core (i.e., one polling thread) for Rx and 6 cores for Tx (while 1 core is reserved for logging/monitoring processes). The Rx thread keeps polling the virtual NIC to fetch the incoming messages and dispatch each message to one Tx thread via the lockless queue, which continues to replicate and forward the messages.
Minimizing packet replication overheads: When a Tx thread is replicating the message, instead of creating packets with each containing one complete copy of the message, we use a zero-copy message replication technique. For each packet, we remove the first few bytes that contain Ethernet and IP header and then invoke API to make several shallow copies of the packet, equal to the number of downstream nodes of a proxy. We allocate small buffers for new Ethernet and IP headers (from pre-allocated memory pools) and attach each pair of these buffers to one shallow copy created previously. Then we configure the headers (i.e., writing the appropriate destination addresses) and use API to send the packets out.
Parallelizing multiple multicast trees: To further improve the throughput, we borrow the sharding idea used by CloudEx [4]. Since each piece of market data is associated with one trading symbol (e.g., $MSFT, $AAPL, $AMD), we can employ multiple trees in parallel to multicast the market data associated with different symbols. In this way, Jasper’s throughput scales horizontally by adding more multicast trees.
Appendix F Receiver Hedging Synchronization
As two VMs belonging to an MP participate in trading simultaneously, (i) their state (e.g., LOB snapshot used for deciding when to place orders) need to be in sync and (ii) we need a way to discard any duplicate order submissions i.e., if both VMs of an MP submit the same order, we need a way to identify and discard one copy.
Order De-duplication: If a receiver VM takes hash of all the market data points it uses to create an order and attaches the hash to the order message then the proxies can easily de-duplicate such orders.
State Synchronization: Most of the messages that originate from the exchnage server and may change some state at the receiver VMs will be seen by both VMs of an MP as packet losses in the outbound side are negligible. Furthermore, the exchange server sequences all its messages so the receiver VMs can easily identify the rare gaps and synchronize with each other to reconcile their state.
Appendix G Appendix: Summary Look
Table 9 provides a quick look at the Jasper’s goals and corresponding techniques.
Goals | Our Techniques |
Multicast | RRPS, Proxy hedging, Receiver hedging, |
Bursty Order Submission | Pacing at proxies, Limit Order Queue |
Security | Trust models |
Appendix H Appendix: Multicast Packet Losses
We define a packet as lost if at least one of multicast receivers do not receive it. In our 20 seconds benchmark on GCP, using c2d-highcpu-8 instances, with 100 multicast receivers, we observe small packet losses: packets lost with a multicast message rate of 350K MPS (utilizing egress of each proxy). At smaller message rates e.g., 100K MPS, we do not observe a single packet loss.
On AWS, using c5.2xlarge instances, we observe slightly greater packet losses: lost packets with a multicast rate of 10K MPS and lost packets with a rate of 100K MPS. Although we are exhausting egress of a proxy at 100K MPS, we suspect that the increased losses are due to the packet per second quotas implemented by AWS [47].